# Nonlocal mass creation by metricized accelerations in Euclidean matterspace

### v1

@@ 4 Shannon information in the yin-yang balance of distributed energies
In Europe, such yin-yang cancellation of kinetic and potential energies for a secluded star or other isolated system was first assumed by Pascual Jordan in 1936, but was not supported quantitatively by relevant calculations. Monistic geometrodynamics, where the flat matterspace of a quasi-isolated system relates the self-generated mass densities to local accelerations in the correlated distribution, returns to the integral balance of kinetic and potential energies under equal active and passive masses. The equivalence of passive (reactive) and active (attractive) mass densities is a mathematical consequence of four covariant relations $T_{o}^{\mu}(x)=0$ for the flat-space continuum of massive fields under the adaptive local time. The vector geometrodynamics of the correlated tensor densities $T_{i}^{j}(x)\neq 0$ in the nonlocal mass-energy organization obey the Bianci identities for relativistic four-accelerations. The latter are the primary reason for the observable existence of the physical world with inertial masses and adaptive changes of their densities.
@@ 5 Conclusion
Local dissipation processes from external (non-metric) interventions or intensive radiation losses of dense elastic regions violate the metric organization with nonlocal correlations in closed mechanical systems. Non-metric energy-momentum exchanges with $\nabla_{\mu}j^{\mu}(x)\neq 0$ are much easier to analyze in the dualistic terms of separated particles in Newton’s mechanical laws, rather than in the Cartesian terms of non-empty space with continuous vortex densities. The dualistic simplification of matter has had tremendous quantitative success in practice, leading to the Standard Model of physics for separated particles and their field mediators.
Despite the philosophical directives and the predicted evolution of physics towards the nondual field theory [15], one can say again that “people slowly accustomed themselves to the idea that the physical state of space itself were finite physical reality”. Cartesian publications on matter-extension and monistic physics are nevertheless still underrepresented in the current volume of dualistic alternatives for mechanical events.
@@
The correlated matter-space or elastic space-plenum of the ancient Greeks can have many unique applications. It consistently accepts the nonlocal nature of elastic self-organization both in the quantum microcosm and in the whole Metagalaxy. Material space-plenum can be continuously filled with bits of information, which cannot be done for emptiness. Therefore, monistic physics has the power to initiate a new informatics of big data based on instantaneous stresses of the elastic cosmos, but not on the slow wave traffic of dissipative exchanges with retardation.
@@
Practical applications of the monistic picture and the geometrodynamics of correlated distributions in macroscopic physics, biology, informatics, cosmology, hydraulic engineering and many other energy-related disciplines go beyond the presented extension of the static self-assembly [12] to its covariant self-motion in the absence of non-metric forces. Finally, the twentieth-century initiatives of Einstein-Grossmann, Mie, and Hilbert should one day return to the flat-space material plenum of the ancient Greeks and the monistic worldview of Descartes. The renaissance of Euclidean subgeometry for material space consistently explains the discovered entanglement of macroscopic states and can provide new testable predictions for further nonlocal experiments.
Data Availability Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest The author has no conflict of interest to declare that are relevant to the content of this article.

### v2

@@ 4.2 Shannon information references nullify energy integral in equilibrium self-assembly
Monistic geometrodynamics, where the flat matterspace of a quasi-isolated system relates the self-generated mass densities to local accelerations in the correlated distribution, returns to the integral cancellation of kinetic and potential self-energies under equal active and passive masses. The equivalence of passive (reactive) and active (attractive) mass densities is a mathematical consequence of four variational equations $T_{o}^{\mu}(x)=0$ for the flat-space continuum of massive fields under the adaptive local time. The vector geometrodynamics of the correlated tensor densities $T_{i}^{\mu}(x)\neq 0$ in the nonlocal mass-energy organization obey the Bianci identities for relativistic four-accelerations. The latter are the primary reason for the observable existence of the physical world with inertial masses and adaptive changes of their densities.
The field monism of living matter, material thoughts and inert or toxic masses in the common energetic and informational filling (13) of Vernadsky’s Geo-Bio-Noosphere does not allow, in principle, to give an independent definition of neither living nor inert (inanimate) matter in a satisfactory way. Due to the gloval superposition in common energy and information content, we can think about the simultaneous emergence of living and inert hierarchies after the Big Bang or the eternal existence of living self-organisations in bouncing models of the Metagalaxy.
5 Experiment can falsify theory, never validate it
5.1 Misinterpretation of world nonlocality in local measurements of energy-momentum exchange
According to Gödel’s theorems, it is impossible to establish a closed algebraic theory without reference to external mathematical constructions. Similarly, there are no completely isolated subsystems in their hierarchical immersion in the physical whole. Nevertheless, for many computational models it is more convenient to consider a quasi-isolated subsystem as an isolated energy body or an elementary particle. For relatively weak mechanical relations between formally isolated masses, such an approximation of reality suggests the field concept of mutual forces between inertial partners. Such historical division of the monistic whole into almost isolated dense regions of the common space of nonlocal mass, energy and information led first to the modeling success of the two-body problem, and then to the triumph of statistical ensembles in dual partitions of the continuous medium into atoms and molecules.
The dualistic worldview not only allowed for the discovery of Neptune at the tip of a pen, but also led to the theoretical standards in simplified physics of massive particles and massless fields modeling long-range interactions. It is believed that since the Standard Model has allowed experiments with highenergy particles to be described with the highest modern precision, monistic alternatives are unable to compete with references to Newtonian mechanics and gravitation on the mega- and macroscale. It is often claimed by Cartesian opponents that since the main predictions of special and general relativity have been quantitatively confirmed in practice, these coherent extensions of Newtonian mechanics and gravitation also preserve the dual concept of localized partners in the ambient interaction field.
On the contrary, Kuhn claimed that “Einstein theory could be accepted only with the recognition that Newton’s was wrong” [21]. Indeed, Einstein’s geodesic motion [10] has four mechanical potentials $g_{\mu}$, while incomplete classical mechanics accepts only one gravitational potential $g_{o}$; masses after warming weigh more only in Einstein’s physics, etc. After decades of explicit theoretical concerns about the accepted interpretation of General Relativity [22, 23], cosmologists still believe in “the well known fact that the weak field, low velocity, low pressure corresponding limit of GR is Newtonian gravity, as evidenced by solar system tests” [24].
The true fact is that the solar system exists in few billions of years, while the dualistic physics of local Newtian masses can explain its mechanical stability only for a few handred years [25]. Instantaneous geometric correlations for self-consistent stresses and mass densities of a nonlocal medium or elastic matterspace in geometrodynamics (7)-(8) can preserve the dynamical self-assembly of the solar system indefinitely if the dissipative exchange with nearby stars is neglected. The spherically symmetric continuum of mass in the non-locally organized matterspace of a secluded star admits not only instantaneous correlations of metric stresses, but also diverging-converging spherical waves of mass density of longitudinal character. Such a cosmic coherence of extended matter around the dense visible densities of the shining star explains three signals in Kozyrev’s telescopic experiments, including the wave echo from the periphery to the center, often interpreted as an ‘advanced’ flux to the future positions of a moving star [26, 27].
After the numerical study of the appearing ‘advanced’ - retarded fronts of converging - diverging longitudinal shifts in the monistic space continuum, one can, for example, predict future locations of earthquakes or weather disasters. Such promising applications of the geometrodynamic equation (7) to the future energy events in the nonlocal Earth continuum or to the medium-term prognosis for the Sun’s activity [28] can be a practical reason for further studies of monistic macrophysics with instantaneous entanglement of stresses and ‘advanced’ returns of bound longitude perturbations or wave ekhos. Geometrodynamics of elastic matterspace can be universally applied to as to megascale self-oscillations in cosmological formations and pulsating stars, as to force-free ‘zitterbewegung’ of microscale of microscale densities [29, 30].
Non-local auto-organization of correlated densities of an isolated system at both scale limits is consistent with self-interaction under conservation of mechanical energy. Such autodynamics does not obey Newtonrqs 2nd law for change of energy-momentum of a considered body by an external force without feedbacks. Local measurements detect apparent energy changes for probe elements, and such dissipative changes are used to misinterpret the conceptual monism of elastic matterspace in dualistic terms of hierarchical partners with individual energy contents.
5.2 Schwartzschild metric with model singularities has not yet been falsified by measurements
Einstein performed a thought experiment in 1939 and logically rejected the Schwarzschild metric solution in the strong field region together with unphysical singularities [14]. The introduction of Mach’s nonlocal ideas into Einstein’s tensor formalism allowed in 2007 to rewrite the 4 -interval into the non-Schwarzschild version [16], $ds^{2}=\left[rdx^{o}/\left(r+r_{o}\right)\right]^{2}-dr^{2}-r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right)$, which quantitatively justified the equivalence of active and passive masses in Euclidean matterspace.
It seems that the December 1915 metric [31] is still forbidden for criticism because it is identified by the mass media as the most famous achievement of Einstein, not Schwarzschild. Theorists of leading gravitational journals try to protect the uniqueness of the Schwarzschild metric by Birkhoff’s theorem. At the same time they ignore the fact that this theorem is valid only for the void and not for nonlocal reality with Cartesian matter extension. Experimentalists are also convinced that ‘the reluctant father of black holes’ was wrong to deny the existence of metric singularities in 1939, since the predictions of the Schwarzschild metric have been very reliably verified by numerous measurements. However, any mathematician can easily check that the 2007 metric for the nonlocal filling of Euclidean space by radial masses with equal passive and active densities leads to the same [32] quantitative predictions as the Schwarzschild modelling of solar fields in well-known GR probes. Remember that measurements do not validate competing theories, but only falsify incorrect approaches as accuracy increase [13]. The strong field predictions from both the 1915 and 2007 4-intervals have not yet been falsified in the available range of measurements.
The second order gravitational redshift can in principle distinguish between curved empty space in dualistic physics and flat matterspace with adaptive time in our monistic alternative. Again, the 3D Euclidean sub-interval for Cartesian matter-extension with $g_{oo}^{2007}=r^{2}/\left(r+r_{o}\right)^{2}$ for the static radial organization maintain the pseudo-Riemannian 4D geometry and the same first-order corrections to Newtonian gravity for Mercury’s perihelion precession, redshift, and light bending as predicted by the Schwarzschild metric with $g_{oo}^{1915}=1-2r_{o}c^{2}/r$ for the main GR tests in the weak negative potential $\tilde{\varphi}=-r_{o}/r$, when $r_{o}\equiv$ $Gm/c^{2}\ll r$. But the second order relativistic corrections are different for the warped empty space and for the Euclidean material space. Precise measurements of the physical time dilation in the Earth-Sun-Moon system with the time-varying weak potential $\tilde{\varphi}$ make it possible to distinguish the 1915 time dilation $(d\tau-dt)/dt\equiv\sqrt{g_{oo}^{1915}}-1\approx\tilde{\varphi}-\left(\tilde{\varphi}^{2}/2\right)$ in curved empty space from alternative time dilation $\sqrt{g_{oo}^{2007}}-1\approx\tilde{\varphi}+\tilde{\varphi}^{2}$ in the flat matterspace.
5.3 3D space flatness from SQUID measurements
So far inhomogeneous gravity and acceleration have never altered SQUID records for macroscopic loops of superconductors. Such null results are consistent with the macroscopic flatness of the laboratory space with the highest quantum accuracy [33]. If anyone were to measure the contribution of gravitational fields or accelerations to the quantization of magnetic flux from self-correlated currents in superconducting contours, the Euclidean geometry of the laboratory space would in principle be falsified. In turn, the experimental independence of the quantized magnetic flux from changes in the gravitational field or changes in the acceleration changes falsifies its theoretical relation to the 3 -space curvature in dualistic interpretations of Einstein’s geometric fields. SQUID-based accelerometers could only exist in a metrically warped 3 -space and have never been built in physical reality. Based on the absence of SQUID accelerometers and gravimeters, new generation of relativists should think about the metricdependent 4-interval with a universal (Euclidean) 3-section and adaptive local time in Einstein’s physics.
5.4 Electrodynamic experiments can falsify the dogma of retarded actions ‘from there to here’
The GR metric forces are definitely not the pulling ‘action-at-a-distance’. They seem to be local geometric pushes, like superpenetrating liquid pushes in Lomonosov’s aetheric model. The pan-unity of visible and invisible matter in the Lomonosov and Umov continuum of kinetic energy with inertial heat means, in modern terms, spatial non-locality of inertial densities with correlated field stresses from General Relativity. By interpreting the mathematical beauty of geometric fields Einstein has publicly admitted the ponderable ether: “… space is endowed with physical qualities; in this sense, therefore, there exists an ether. According to the general theory of relativity space without ether is unthinkable; for in such space there not only would be no propagation of light, but also no possibility of existence for standards of space and time (measuring-rods and clocks), nor therefore any space-time intervals in the physical sense. But this ether may not be thought of as endowed with the quality characteristic of ponderable media, as consisting of parts which may be tracked through time. The idea of motion may not be applied to it” [7].
Time-varying stresses are simultaneous in the nonlocal field theory, maintaining constant energy integrals at all world moments $x^{o}$. Wave modulations of correlated densities can be retarded (diverging from the mass vertex) or ‘advanced’ (converging to it). The monistic GR formalism for continuous matterspace with dense (visible) and rarefied (invisible) regions may shed some light on the large-scale correlation of galaxy motions [9], the megaparsec alignment of the pulsars [8], and the macroscopic entanglement of photons in the Earth experiments commented on for the 2022 Nobel Prize in Physics.
The instantaneous action on the telescope detector [26, 27] under the actual direction to a moving star suggests timeless Newton and Coulomb forces without temporal retardation due to the intuitive ‘speed of gravity’, estimated by Laplacce above $6\times 10^{6}c$ [25]. The Coulomb’s forces have indeed exhibited the rigid motion with high-energy beams of ultrarelativistic electrons [34]. These well-executed experiments at the Frascati National Laboratory are consistent in principle with the monistic concept of nonlocal matterspace. The modified scheme with high-enegy beams of relativistic charges was also proposed [35] to falsify the empty space model beyong reasonable doubts. A clear rejection of conceptual exprements can provide for the competing worldviews of monists and dualists. The non-local reading of continuous electric charges and mechanical masses supports the local self-action mechanism for the inverse square forces [36]. In such monistic physics, correlated geodesic accelerations push inertial densities ‘from here to there’ instead of the classical pulling ‘from there’, as Newtonian action-at-a-distance states.
Since local self-accelerations define the long-range interactions in the nonlocal self-assembly of matterspace, the intervention of external forces in local regions of such self-assembly resonating with auto-oscillations can perturb the correlated dynamics of distant densities. Non-metric external forces perturb the time-averaged equilibrium of the nonlocal system with the metric energy-information potential $\mathrm{W}$, which in (13) leads to the mutual compensation of the kinetic (yang) and metric (yin) energy integrals. Such a spooky action at a distance by a non-locally organised system was originally predicted from the paradoxes of quantum mechanics with instantaneous reduction of the wave function [37]. Encountered numbers of non-local laboratory experiments and spooky communications between bio-cells, bacteria, animals and humans are outside the immediate comments of monistic field physics for macro-correlated hierarchies of Shannon information in geometrized mass-energy.
@@ 6 Conclusion
Observations of the correlation of cosmic matter on megaparsecs in the tenth decade and the Nobel Committee Prize for macroscopic experiments violating Bell’s inequality in 2022 have ended conceptual disputes about the exit of quantum nonlocality from the notorious microcosm to all spatial scales. Renewed interest in the monistic Cartesian matter-extension and field theories of continuous mass distribution has prompted a covariant generalisation of previously geometrized matter for static equilibrium [16].
The formal division of a highly inhomogeneous medium into localised carriers and virtually empty space has led to very successful mechanical and electrical models with remote interaction between the designated partners. The calculation of model thrusts ‘from there to here’ for massive points using the geometric law of inverse squares is much easier than the estimation of non-local self-accelerations ‘from here to there’ in a non-local medium with timeless stresses [36]. The simplified modelling always has its price - the empty space without elasticity cannot lead to the time-averaged equilibrium of self-gravitating systems and gives rise to many puzzles in astronomical observations, starting with the Bentley’s paradox.
@@
Local dissipation processes from external (non-metric) interventions or intensive radiation losses of dense elastic regions violate the metric organization with nonlocal correlations in closed mechanical systems. Non-metric energy-momentum exchanges with $\nabla_{\mu}j^{\mu}(x)\neq 0$ are much easier to analyze in the dissipative terms of the Newton Second Law, rather than in the Cartesian terms of integral energy-momentum exchanges with continuous vortex densities. The dualistic reinterpretation of purely field matter has had tremendous quantitative success for practice, leading to the Standard Model of physics for separated particles and their field mediators. Despite the philosophical directives and the predicted evolution of physics towards the nondual field theory [15], one can say again that “people slowly accustomed themselves to the idea that the physical states of space itself were the finite physical reality”. Cartesian publications on matter-extension and monistic physics are still underrepresented in the current volume of dualistic alternatives for mechanical events.
@@
The correlated matter-space or elastic space-plenum of the ancient Greeks can have many unique applications. It consistently accepts the nonlocal nature of elastic self-organization both in the quantum microcosm and in the whole Metagalaxy. Material space-plenum can be continuously filled with bits of information, which cannot be done for emptiness. Therefore, monistic physics has the power to initiate a new informatics of big data based on instantaneous stresses of the elastic cosmos, but not only on the slow wave traffic of dissipative exchanges with retardation.
@@
Practical applications of the monistic picture and the geometrodynamics of correlated distributions in macroscopic physics, biology, informatics, cosmology, hydraulic engineering and many other energy-related disciplines go beyond the presented extension of the static self-assembly to its covariant self-motion in the absence of non-metric forces. Finally, the twentieth-century initiatives of Einstein-Grossmann and Mie should one day reinforce the flat matter-space-plenum of the ancient Greeks, the vortex matter-extension of Descartes, and the etheric world pan-unity in the monistic terms of Russian Cosmism. The renaissance of Euclidean 3-space for correlated inertial fields addresses the discovered entanglement of macroscopic states and bridges relativistic physics with quantisation, advancing geometrodynamics to decipher the non-local transfer of big data in the material field reality.
Acknowledgments. The author is grateful to Prof. Guido Pizzella for visiting Frascati National Laboratories to study the details of the conceptual experiment [34].